# Preparation of Papers for IEEE Sponsored Conferences & Symposia*

### v1

@@ V-C1 Permitted Execution Times
Since constraints under our branching rule restrict move execution and vertex occupancy over time intervals (Definition V.1), every set $\mathcal{T}_{x}$ where $x$ ranges over all moves $m\in\mathcal{A}$ and vertices $v\in\mathcal{V}$ can be described as the union of a finite number maximally connected time intervals:
$\mathcal{T}_{x}=\bigcup_{h=1}^{N_{x}}S_{x}^{h},\quad S_{x}^{h}=[s_{x}^{h},e_{x}^{h})$ (11)
@@
V-C2 CSIPP Planning Behavior
@@
CSIPP performs an A [30] search in the vertex-safe interval space. Each search state is represented by a tuple $(v,S_{v})$ where agent $i$ is located at $v\in\mathcal{V}$ during the safe interval $S_{v}=[s_{v},e_{v})\subseteq\mathcal{T}_{v}$. Let $t^{a}\in S_{v}$ be the time when $i$ arrives $v$. Unless state $(v,S_{v})$ is a goal state and $e_{v}=\infty$, agent $i$ must execute a move action to leave $v$ before the end of the safe interval at $e_{v}$. For each move action $m$ with $\mathit{from}(m)=v$, agent $i$ can execute $m$ at any time $t$ satisfying
  • $t\in[t^{a},e_{v})$ (after arriving and before it is unsafe),
  • $t\in\mathcal{T}_{m}$ (it is safe to execute the move action), and
  • $t+m_{D}\in\mathcal{T}_{v^{\prime}}$ (it is safe to occupancy $v^{\prime}$),
where $v^{\prime}=\mathit{to}(m)$. In other words, $t\in P_{m}$ where
@@
and $\mathcal{T}_{v^{\prime}}^{m_{D}}=\{t-m_{D}\mid t\in\mathcal{T}_{v^{\prime}}\}$ to compensate for the arrival at $v^{\prime}$ being $m_{D}$ after $t$. By (11), $\mathcal{T}_{m}$ and $\mathcal{T}_{v^{\prime}}^{m_{D}}$ are unions of maximally connected intervals, such that $P$ is also a union of maximally connected intervals. Fig. 8 illustrates how $P$ is formed from the intersection between $t^{a},e_{v})$, $\mathcal{T}_{m}$ and $\mathcal{T}_{v^{\prime}}^{m_{D}}$.
@@
Figure 8: Illustration of the construction of $P$ as the intersection between $t^{a},e_{v})$, $\mathcal{T}_{m}$ and $\mathcal{T}_{v^{\prime}}^{m_{D}}$.
@@
For every maximally connected interval $s_{m},e_{m})\subseteq P_{m}$, an edge representing $i$ executing $\langle m,s_{m}\rangle$ is created, leading to the next state $(v^{\prime},S_{v^{\prime}})$ where $S_{v^{\prime}}\subseteq\mathcal{T}_{v^{\prime}}$ is the safe interval at $v^{\prime}$ such that $s_{m}+m_{D}\in S_{v^{\prime}}$.
@@
CSIPP always selects the earliest time $s_{m}$ for $i$ to execute $m$. This earliest-time selection property is essential for our proof: since the earliest time $s_{m}$ in $s_{m},e_{m})$ is selected, it is not possible to select an earlier time $t^{\prime}<s_{m}$ arbitrarily close to $s_{m}$ for the move action $m$ to be executed.
V-C3 Non-degenerate Reductions For Move Actions
@@
When our branching rule is applied to a move action $\langle m^{i},t^{i}\rangle\in\pi_{i}$ in a move-move or move-wait conflict, a constraint is created which forbids the execution of $m^{i}$ during a non-degenerate interval $t^{i},t^{\prime})$. We assume that $\pi_{i}$ was returned from CSIPP, such that a safe interval $s_{m},e_{m})\subseteq P_{m}$ with $s_{m}=t_{i}$ was identified during the search. Consequently, the new constraint removes either a prefix or the entirety of $s_{m},e_{m})$ from $P_{m}$.
@@
Starting with a move-move conflict, our branching rule forbids the execution of $m^{i}$ during the unsafe interval $t^{i},t^{i}_{u})$, which is non-degenerate by construction.
@@
Consider instead a move-wait conflict $\langle\langle m^{i},t^{i}\rangle,\langle w^{j},t^{j}\rangle\rangle$ with collision interval $I^{c}=[t_{1}^{c},t_{2}^{c})$ and intersection interval $\bar{I^{c}}=[\bar{t_{1}^{c}},\bar{t_{2}^{c}})$. Our branching rule creates a constraint forbidding the execution of $m^{i}$ during $t^{i},t^{i}+\delta)$. If $\delta$ is bounded away from 0, then $t^{i},t^{i}+\delta)$ is non-degenerate. By (7), $\delta=\min\left(\gamma|\bar{I^{c}}|,t^{j}+w^{j}_{D}-\bar{t_{1}^{c}}\right)$ is the minimum of two positive quantities:
@@
  • $\delta=\gamma|\bar{I^{c}}|$: It holds that $\delta$ is bounded away from zero since $\bar{I^{c}}$ is non-degenerate and $\gamma\in(0,1)$ is a fixed, non-zero value.
@@
  • $\delta=t^{j}+w^{j}_{D}-\bar{t_{1}^{c}}$: From (5) we know that $I^{c}\subseteq\bar{I^{c}}$ and $I^{c}\subseteq[t^{j},t^{j}+w^{j}_{D}]$. The former implies that $\bar{t_{1}^{c}}\leq t_{1}^{c}<t_{2}^{c}\leq\bar{t_{2}^{c}}$ where strict inequality holds since $I^{c}$ is non-degenerate. The latter implies that $t^{j}\leq t_{1}^{c}<t_{2}^{c}<{t_{2}^{c}}^{\prime}\leq t^{j}+t^{j}_{D}$ where ${t_{2}^{c}}^{\prime}>t_{2}^{c}$ is arbitrarily close to $t_{2}^{c}$ (to correct for $I^{c}=[t_{1}^{c},t_{2}^{c})$ not including $t_{2}^{c}$ while $t^{j},t^{j}+t^{j}_{D}]$ does include $t^{j}+t^{j}_{D}$). These together give
@@
    $0$ $<t^{j}+w^{j}_{D}-\bar{t_{1}^{c}}=\delta,$
    such that $\delta$ is bounded away from 0.
Hence, $\delta$ is bounded away from zero, implying that the constrained interval $t^{i},t^{i}+\delta)$ is non-degenerate.
In both cases, where $\langle m^{i},t^{i}\rangle\in\pi_{i}$ is contained in a move-move or move-wait conflict, a constraint is created forbidding the execution of $m^{i}$ during a non-degenerate interval $t^{i},t^{\prime})$. That is, $t^{i},t^{\prime})$ is removed from $\mathcal{T}_{m}$. Since $s_{m}=t^{i}$ for some $s_{m},e_{m})\subseteq P_{m}$, by (12), either the entire safe interval $s_{m},e_{m})$ is removed (if $t^{\prime}\geq e_{m}$) or a non-degenerate prefix $s_{m},t^{\prime})$ is removed from $P_{m}$ (if $t^{\prime}<e_{m}$).
V-C4 Non-degenerate Reductions For Wait Actions
In this section, we consider the constraints created when applying our branching rule to a wait action $\langle w^{j},t^{j}\rangle\in\pi_{j}$ in a move-wait conflict $\langle\langle m^{i},t^{i}\rangle,\langle w^{j},t^{j}\rangle\rangle$. It will be shown that this leads to two cases. In the first case, we show that there exists a timed move action $\langle m,t^{j}+w^{j}_{D}\rangle\in\pi_{j}$ following $\langle w^{j},t^{j}\rangle$, and a constraint is created that forbids the execution of $m$ during a non-degenerate interval $t^{j}+w^{j}_{D},t^{\prime})$. Therefore, the same conclusions as in Section V-C3 are drawn. In the second case, a constraint is created forbidding the use of $v(w^{j})$ during a non-degenerate interval $t^{\prime},t^{\prime\prime})$, necessarily leading to a non-degenerate reduction in $\mathcal{T}_{v(w^{j})}$.
When our branching rule is applied to $\langle w^{j},t^{j}\rangle$, two constraints are created (Definition V.1). One constraint forbids $j$ from occupying $v(w^{j})$ during $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})$, and the other forbids the execution of any move action $m$ with $\mathit{from}(m)=v(w^{j})$ at any time in $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})$. We begin by showing that the constrained interval $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})$ is non-degenerate by expanding $\bar{t_{1}^{c}}+\delta$ using (7):
@@ V-C4 Non-degenerate Reductions For Wait Actions
$=\min\left((1-\gamma)\bar{t_{1}^{c}}+\gamma\bar{t_{2}^{c}},\;t^{j}+w^{j}_{D}\right)$ (13)
@@
which shows that $\bar{t_{1}^{c}}+\delta\leq(1-\gamma)\bar{t_{1}^{c}}+\gamma\bar{t_{2}^{c}}$. Since $0<\gamma<1$ is fixed to some value less than $1$, it follows that $(1-\gamma)\bar{t_{1}^{c}}+\gamma\bar{t_{2}^{c}}<\bar{t_{2}^{c}}$ and therefore that $\bar{t_{1}^{c}}+\delta<\bar{t_{2}^{c}}$. Thus, $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})$ is non-degenerate.
Next, we show that the constrained $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})$ necessarily overlap with the time interval $t^{j},t^{j}+w^{j}_{D}]$ when $j$ occupies vertex $v(w^{j})$ while executing $\langle w^{j},t^{j}\rangle$. To do so, we combine two conditions. First, under the assumption made in Section III-B that $I^{c}$ is non-degenerate together with $I^{c}\subseteq[t^{j},t^{j}+w^{j}_{D}]$ by (5), it follows that $t^{j}<\bar{t_{2}^{c}}$ meaning that $j$’s occupancy of $v(w^{j})$ starts before the constrained interval ends. Second, (13) shows that $\bar{t_{1}^{c}}+\delta\leq t^{j}+w^{j}_{D}$, implying that $j$’s occupancy of $v(w^{j})$ ends at or after the constrained interval begins. Together, these conditions are sufficient to guarantee overlap by at least a singular value.
There exists two cases: the overlap between $t^{j},t^{j}+w^{j}_{D}]$ and $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})$ is either singular or non-degenerate.
@@
  • Singular: this implies that $\bar{t_{1}^{c}}+\delta=t^{j}+w^{j}_{D}$. In this case, $w^{j}$ cannot be the final infinite wait action in $j$’s plan: $\bar{t_{1}^{c}}<\infty$ (since collisions occur at finite times) and $\delta<\infty$ (since $|\bar{I^{c}}|<\infty$, see (7)). Thus, the left-hand-side is $\infty$ implying that the right-hand side containing $w^{j}_{D}$ is also $\infty$. Since $w^{j}$ does not have infinite duration and is therefore not the final infinite wait action, there is a subsequent move action $\langle m,t^{j}+w^{j}_{D}\rangle\in\pi_{j}$ with $\mathit{from}(m)=v(w^{j})$ in $j$’s plan. Our branching rule creates a constraint forbidding $m$ from being executed during the non-degenerate interval $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})=[t^{j}+w^{j}_{D},\bar{t_{2}^{c}})$. The same reasoning and conclusions as in Section V-C3 for move actions in move-move and move-wait conflicts applies: a non-degenerate prefix or the entirety of some safe interval $s_{m},e_{m})\subseteq P_{m}$ with $s_{m}=t^{j}+w^{j}_{D}$ identified during the CSIPP search is removed from $P_{m}$.
@@
  • Non-degenerate: we know that $t^{j},t^{j}+w^{j}_{D}]\subseteq\mathcal{T}_{v(w^{j})}$ else the timed wait action $\langle w^{j},t^{j}\rangle$ would not have been returned by CSIPP. This means that the non-degenerate overlap $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})\cap[t_{j},t^{j}+w^{j}_{D}]$ is also a subset of $\mathcal{T}_{v(w^{j})}$. In other words, this tells us that a non-degenerate part of $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})$ exists in $\mathcal{T}_{v(w^{j})}$. Our branching rule creates a constraint forbidding the occupancy of $v(w^{j})$ during $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})$, therefore removing $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})$ from $\mathcal{T}_{v(w^{j})}$. Thus, a non-degenerate interval is removed from $\mathcal{T}_{v(w^{j})}$.
V-C5 Bounded Plan Impossibility

### v2

@@ 5.3.1 Permitted Execution Times
Since constraints under the various branching rules restrict move execution and vertex occupancy over time intervals (Definition 5.1), every set $\mathcal{T}_{x}$ — with $x$ ranging over all moves $m\in\mathcal{A}$ and vertices $v\in\mathcal{V}$ — can be described as the union of a finite number of maximally connected time intervals:
$\mathcal{T}_{x}=\bigcup_{h=1}^{N_{x}}S_{x}^{h},\quad S_{x}^{h}=[s_{x}^{h},e_{x}^{h}).$ (11)
@@
5.3.2 CSIPP Planning Behavior
@@
CSIPP performs an A [31] search in the vertex-safe interval space. Each search state is represented by a tuple $(v,S_{v})$ where agent $i$ is located at $v\in\mathcal{V}$ during the safe interval $S_{v}=[s_{v},e_{v})\subseteq\mathcal{T}_{v}$. Let $t^{a}\in S_{v}$ be the time when $i$ arrives at $v$. Unless state $(v,S_{v})$ is a goal state and $e_{v}=\infty$, agent $i$ must execute a move action to leave $v$ before the end of the safe interval, $e_{v}$. For each move action $m$ with $\mathit{from}(m)=v$, agent $i$ can execute $m$ at any time $t$ satisfying
  • $t\in[t^{a},e_{v})$ (executing $m$ after arriving at $v$ and before it is unsafe to remain there),
  • $t\in\mathcal{T}_{m}$ (executing $m$ when it safe to do so), and
  • $t+m_{D}\in\mathcal{T}_{v^{\prime}}$ (executing $m$ to arrive at the next node $v^{\prime}=\mathit{to}(m)$ when it is safe to do so),
In other words, it is safe to execute $m$ at some time $t\in P_{m}$ where
@@
and $\mathcal{T}_{v^{\prime}}^{m_{D}}=\{t-m_{D}\mid t\in\mathcal{T}_{v^{\prime}}\}$ to compensate for the arrival at $v^{\prime}$ being $m_{D}$ after $t$. By (11), $\mathcal{T}_{m}$ and $\mathcal{T}_{v^{\prime}}^{m_{D}}$ are unions of maximally connected intervals, such that $P_{m}$ is also a union of maximally connected intervals. Figure 8 illustrates how $P_{m}$ is formed from the intersection between $t^{a},e_{v})$, $\mathcal{T}_{m}$ and $\mathcal{T}_{v^{\prime}}^{m_{D}}$.
@@
Figure 8: Illustration of the construction of $P_{m}$ as the intersection between $t^{a},e_{v})\subseteq\mathcal{T}_{v}$, $\mathcal{T}_{m}$, and $\mathcal{T}_{v^{\prime}}^{m_{D}}$.
@@
For every maximally connected interval $s_{m},e_{m})\subseteq P_{m}$, CSIPP creates an edge representing that agent $i$ executes $\langle m,s_{m}\rangle$, leading to the next state $(v^{\prime},S_{v^{\prime}})$ where $S_{v^{\prime}}\subseteq\mathcal{T}_{v^{\prime}}$ is the safe interval at $v^{\prime}$ such that $s_{m}+m_{D}\in S_{v^{\prime}}$.
@@
CSIPP always selects the earliest time $s_{m}$ in a safe interval $s_{m},e_{m})\subseteq P_{m}$ for $i$ to execute $m$. This earliest-time selection property is essential for our proof: since the earliest time $s_{m}$ in $s_{m},e_{m})$ is selected, it is not possible to select an earlier time $t^{\prime}<s_{m}$ arbitrarily close to $s_{m}$ for the move action $m$ to be executed, since $t^{\prime}\not\in P_{m}$.
5.3.3 Non-degenerate Reductions For Move Actions
@@
When $\delta$-BR is applied to a move action $\langle m^{i},t^{i}\rangle\in\pi_{i}$ in a move-move or move-wait conflict, the motion constraint $\langle i,m^{i},[t^{i},t^{\prime})\rangle$ is created. Assuming that $\pi_{i}$ was returned from CSIPP, a safe interval $s_{m^{i}},e_{m^{i}})\subseteq P_{m^{i}}$ with $s_{m^{i}}=t_{i}$ was identified during the search to construct $\pi_{i}$. We will first show that the constrained interval $t^{i},t^{\prime})$ is non-degenerate, and then consequently that the new constraint removes either a prefix or the entirety of $s_{m^{i}},e_{m^{i}})$ from $P_{m^{i}}$.
@@
If $\langle m^{i},t^{i}\rangle$ is contained in a move-move conflict, $\delta$-BR forbids the execution of $m^{i}$ during the unsafe interval $t^{i},t^{i}_{u})$ which is non-degenerate by construction. Thus, in this case we know that the constrained interval $t^{i},t^{\prime})=[t^{i},t^{i}_{u})$ is non-degenerate.
@@
Consider instead that $\langle m^{i},t^{i}\rangle$ is contained in a move-wait conflict $\langle\langle m^{i},t^{i}\rangle,\langle w^{j},t^{j}\rangle\rangle$ with intersection interval $\bar{I^{c}}=[\bar{t_{1}^{c}},\bar{t_{2}^{c}})$ and collision interval $I^{c}=[t_{1}^{c},t_{2}^{c})$. $\delta$-BR creates the motion constraint $\langle i,m^{i},[t^{i},t^{i}+\delta)\rangle$. The constrained interval $t^{i},t^{\prime})=[t^{i},t^{i}+\delta)$ is non-degenerate only if $\delta$ is bounded away from 0. By (7), $\delta$ is the minimum of two positive quantities, $\gamma|\bar{I^{c}}|$ and $t^{j}+w^{j}_{D}-\bar{t_{1}^{c}}$:
@@
  • $\gamma|\bar{I^{c}}|$ is bounded away from zero since $\bar{I^{c}}$ is non-degenerate and $\gamma\in(0,1)$ is a fixed, non-zero value.
@@
  • From (5) we know that $I^{c}\subseteq\bar{I^{c}}$ and $I^{c}\subseteq[t^{j},t^{j}+w^{j}_{D}]$. The former implies that $\bar{t_{1}^{c}}\leq t_{1}^{c}<t_{2}^{c}\leq\bar{t_{2}^{c}}$ where strict inequality holds since $I^{c}$ is non-degenerate. The latter implies that $t^{j}\leq t_{1}^{c}<t_{2}^{c}<{t_{2}^{c}}^{\prime}\leq t^{j}+t^{j}_{D}$ where ${t_{2}^{c}}^{\prime}>t_{2}^{c}$ is arbitrarily close to $t_{2}^{c}$ (to correct for $I^{c}=[t_{1}^{c},t_{2}^{c})$ not including $t_{2}^{c}$ while $t^{j},t^{j}+t^{j}_{D}]$ does include $t^{j}+t^{j}_{D}$). These together give
@@
    $0$ $<t^{j}+w^{j}_{D}-\bar{t_{1}^{c}}.$
    Thus, $t^{j}+w^{j}_{D}-\bar{t_{1}^{c}}$ is bounded away from zero.
Since $\delta$ is the minimum of two positive quantities which are both bounded away from zero, $\delta$ is also bounded away from zero. Therefore, $t^{i},t^{\prime})=[t^{i},t^{i}+\delta)$ is non-degenerate.
In both cases — when $\langle m^{i},t^{i}\rangle$ is contained in a move-move or move-wait conflict — a constraint $\langle i,m^{i},[t^{i},t^{\prime})\rangle$ is created for which we have shown above that $t^{i},t^{\prime})$ is non-degenerate. This constraint effectively removes $t^{i},t^{\prime})$ from $\mathcal{T}_{m^{i}}$. Since $t^{i}=s_{m^{i}}$ for some $s_{m^{i}},e_{m^{i}})\subseteq P_{m^{i}}$, by (12), either the entire safe interval $s_{m^{i}},e_{m^{i}})$ is removed (if $t^{\prime}\geq e_{m^{i}}$) or a non-degenerate prefix $s_{m^{i}},t^{\prime})$ is removed from $P_{m^{i}}$ (if $t^{\prime}<e_{m^{i}}$). This is a useful result as it shows that $\delta$-BR applied to any timed move action in a conflict necessarily leads to a non-degenerate time interval being removed from the possible times to execute the move action.
5.3.4 Non-degenerate Reductions For Wait Actions
In this section, we consider the constraints created by $\delta$-BR when applied to a wait action $\langle w^{j},t^{j}\rangle\in\pi_{j}$ in a move-wait conflict $\langle\langle m^{i},t^{i}\rangle,\langle w^{j},t^{j}\rangle\rangle$ with intersection interval $\bar{I^{c}}=[\bar{t_{1}^{c}},\bar{t_{2}^{c}})$ and collision interval $I^{c}=[t_{1}^{c},t_{2}^{c})$. $\delta$-BR creates the vertex constraint $\langle j,v(w^{j}),[\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})\rangle$ and motion constraints $\langle j,m,[\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})\rangle$ for each $m\in\mathcal{A}:\mathit{from}(m)=v(w^{j})$. It will be shown that there exist two cases: when the overlap between the constrained interval $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})$ and the wait interval $t^{j},t^{j}+w_{D}^{j}]$ is either singular or non-degenerate. For each case, we will show that either the same conclusions as in Section 5.3.3 can be drawn for the move action $\langle m,t^{j}+w^{j}_{D}\rangle$ following $\langle w^{j},t^{j}\rangle$ in $\pi_{j}$, or a non-degenerate reduction occurs in $\mathcal{T}_{v(w^{j})}$.
We begin by showing that the constrained interval $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})$ is non-degenerate by expanding $\bar{t_{1}^{c}}+\delta$ using (7):
@@ 5.3.4 Non-degenerate Reductions For Wait Actions
$=\min\left((1-\gamma)\bar{t_{1}^{c}}+\gamma\bar{t_{2}^{c}},\;t^{j}+w^{j}_{D}\right).$ (13)
@@
This shows that $\bar{t_{1}^{c}}+\delta\leq(1-\gamma)\bar{t_{1}^{c}}+\gamma\bar{t_{2}^{c}}$. Since $0<\gamma<1$ is fixed to some value less than $1$, it follows that $(1-\gamma)\bar{t_{1}^{c}}+\gamma\bar{t_{2}^{c}}<\bar{t_{2}^{c}}$ and therefore that $\bar{t_{1}^{c}}+\delta<\bar{t_{2}^{c}}$. Thus, $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})$ is non-degenerate.
Next, we show that $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})$ necessarily overlaps with the wait interval $t^{j},t^{j}+w^{j}_{D}]$. To do so, we combine two conditions. First, $I^{c}$ is assumed in Section 3.2 to be non-degenerate, and by (5) we have $I^{c}\subseteq[t^{j},t^{j}+w^{j}_{D}]$. Thus, it follows that
$t^{j}<\bar{t_{2}^{c}}.$ (14)
Second, (13) shows that $\bar{t_{1}^{c}}+\delta\leq t^{j}+w^{j}_{D}$. These two conditions — $t^{j}<\bar{t_{2}^{c}}$ and $\bar{t_{1}^{c}}+\delta\leq t^{j}+w^{j}_{D}$ — together are sufficient to guarantee overlap between $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})$ and $t^{j},t^{j}+w^{j}_{D}]$ by at least a singular value. We now consider each of the cases of singular or non-degenerate overlap:
@@
  • Singular: This implies that $\bar{t_{1}^{c}}+\delta=t^{j}+w^{j}_{D}$. In this case, $w^{j}$ cannot be the final infinite wait action in $j$’s plan: $\bar{t_{1}^{c}}<\infty$ (since collisions occur at finite times) and $\delta<\infty$ (since $|\bar{I^{c}}|<\infty$, see (7)). Thus, the left-hand-side is $\infty$ implying that the right-hand side containing $w^{j}_{D}$ is also $\infty$. Since $w^{j}$ does not have infinite duration and is therefore not the final infinite wait action, there exists a subsequent move action $\langle m,t^{j}+w^{j}_{D}\rangle\in\pi_{j}$ with $\mathit{from}(m)=v(w^{j})$. $\delta$-BR creates a constraint forbidding $m$ from being executed during the non-degenerate interval $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})=[t^{j}+w^{j}_{D},\bar{t_{2}^{c}})$. The same reasoning and conclusions as in Section 5.3.3 for move actions in move-move and move-wait conflicts applies: a non-degenerate prefix or the entirety of some safe interval $s_{m},e_{m})\subseteq P_{m}$ with $s_{m}=t^{j}+w^{j}_{D}$ identified during the CSIPP search is removed from $P_{m}$.
@@
  • Non-degenerate: We know that $t^{j},t^{j}+w^{j}_{D}]\subseteq\mathcal{T}_{v(w^{j})}$ else the timed wait action $\langle w^{j},t^{j}\rangle$ would not have been returned by CSIPP. This means that the non-degenerate overlap $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})\cap[t_{j},t^{j}+w^{j}_{D}]$ is also a subset of $\mathcal{T}_{v(w^{j})}$. In other words, this tells us that a non-degenerate part of $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})$ exists in $\mathcal{T}_{v(w^{j})}$. $\delta$-BR creates a vertex constraint $\langle j,v(w^{j}),[\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})\rangle$ and thereby removes $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})$ from $\mathcal{T}_{v(w^{j})}$. Thus, a non-degenerate interval is removed from $\mathcal{T}_{v(w^{j})}$.
5.3.5 Bounded Plan Impossibility